Find the missing values that solve this equation.
2x+y = 6
(2,____) (0,____) (1,____)

Answer by funmath(2925)

Find the missing values that solve this equation.
2x+y = 6
(2,____) (0,____) (1,____)
:
Ordered pairs are (x,y)
:
For (2,_), let x be 2 and solve for y:
2(2)+y=6
4+y=6
-4+4+y=6-4
y=2
The ordered pair is: (2,2)
:
For (0,_), let x = 0 and solve for y:
2(0)+y=6
0+y=6
y=6
The ordered pair is: (0,6)
:
For (1,_), let x=1 and solve for y:
2(1)+y=6
2+y=6
-2+2+y=6-2
0+y=4
y=4
The ordered pair is (1,4)
Happy Calculating!!!